• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:03:29 on December19,2003

    Entry number 114628

    keyword=Owl shift summary

    Shift Workers

    Peter Monaghan
    Alexandre Camsonne
    Luigi Lagamba

    00:00 Shift begins with checking the steering of the beam with right
    Septum magnet. Unfortunately, the dump ion chambers tripped with the
    septum current at about 88A. We have turned the septum off and we will
    not use it tonight.

    00:40 Put in BeO target; check the beam spot size and also watch
    the Septum temperatures. Everything looks fine - see separate Halog

    00:45 Moved to the thick carbon-12 target. Now we are going to ramp
    up the beam current from 5uA to 100uA and check the septum heating.
    Note, the raster was OFF for this test. Results are shown in the strip charts in a separate Halog entry.

    02:25 Finished the septum heating study. One point of interest was
    that the dump ion chambers tripped when MCC were sending a beam of
    around 95 uA. The change in temperature for each septum (lower/upper coils) was (from 5 - 100uA), right - 0.5/0.45 K and left - 0.3/0.4 K.

    02:30 Now going to do a Bullseye scan with 5uA unrastered beam. Harp scan will be done first for each point before the Coda is started.

    06:30 MCC having optics problems which they are working on to fix.

    06:40 MCC have fixed their optics problem; now setting up for the
    penultimate point on the Bullseye scan.

    07:00 Bullseye scan finally complete; each run was taking a long
    time to register the required number of events since both septum
    magnets and the left HRS were turned off. The last two runs had only
    5000 events due to time constraints to have the hall brought up to
    Restricted access before 8am.

    07:00 Asked MCC to bring Hall A to Restricted Access so that it
    would be completed before 8am.

    07:05 Empty target moved into place.

    07:55 Hall now in restricted access.

    Run Summary

    Run Number Harp Scan run number Comments

    1414 N/A Cosmic data run
    1415 1914 Bullseye scan (0,0) (-238,373)
    1416 1915 Bullseye scan (1,0) (-1226,365)
    1417 1916 Bullseye scan (1,1) (-1206,1427)
    1418 1917 Bullseye scan (0,1) (-211,1419)
    1419 1918 Bullseye scan (-1,1) (719,1370)
    1420 1919 Bullseye scan (-1,0) (682,369)
    1421 1920 Bullseye scan (-1,-1) (650,-625)
    1422 1921 Bullseye scan (0,-1) (-306,-547)
    1423 1922 Bullseye scan (1,-1) (-1233,-532)
    1424 1923 Bullseye scan (0,0) (-257,392)